# Bilayer Excitons in the Laughlin Fractional Quantum Hall State

### v2

@@
thanks: These authors contributed equally to this work.thanks: These authors contributed equally to this work.thanks: These authors contributed equally to this work.thanks: Email: jia__\__li@brown.edu
@@
Ron Q. Nguyen1    Naiyuan J. Zhang1    Navketan Khurana-Batra1,2    Xiaoxue Liu3    Kenji Watanabe4    Takashi Taniguchi5    D. E. Feldman1,2    J.I.A. Li1 1Department of Physics, Brown University, Providence, Rhode Island 02912, USA 2Brown Theoretical Physics Center, Brown University, Providence, Rhode Island 02912, USA 3Tsung-Dao Lee Institute and School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China 4Research Center for Electronic and Optical Materials, National Institute for Materials Science, 1-1 Namiki, Tsukuba 305-0044, Japan 5Research Center for Materials Nanoarchitectonics, National Institute for Materials Science, 1-1 Namiki, Tsukuba 305-0044, Japan
(July 5, 2025)
@@
Figure 1: Exciton pairing between a pair of Laughlin states. (a) Schematic of the Laughlin state in a monolayer, stabilized by intralayer Coulomb interaction. At $1/3$ LL filling, quasiparticle excitations of the Laughlin state carry $1/3$ of electronic charge. (b) Schematic of two Laughlin states in a quantum Hall bilayer, each stabilized by intralayer Coulomb interaction. In this situation, quasiparticles and quasiholes interact through weak interlayer correlation. (c) Schematic diagram showing FQHE within the CF model, where the attachment of two flux quanta is included in the CF construction  Jain.03 . Blue and red solid lines denote integer values for $\nu^{\ast}_{1}$ and $\nu^{\ast}_{2}$, respectively  Li2019pairing ; Liu2019interlayer . At the intersection of these lines, both layers simultaneously develop a robust charge gap, leading to an incompressible state across the quantum Hall bilayer. (d-e) Transport response along $\nu_{1}^{\ast}=1$, at (d) $B=12$ T and (e) $B=28$ T. The top, middle, and bottom panels show PF conductance $G_{\text{PF}}$, counterflow conductance $G_{\text{counter}}$, and drag ratio $I_{\text{drag}}/I_{\text{drive}}$, respectively. Inset of each panel shows schematic measurement configurations in the parallel flow, counterflow, and drag geometries. Blue, green, and red circles mark different ground state orders.
Figure 2: Pairing between quasiparticle excitations. (a,b) Schematic diagrams illustrating (a) thermally excited quasiparticles and quasiholes across the charge gap of the Laughlin state, and (b) suppressed quasiparticle excitation when the charge gap greatly exceeds $k_{B}T$. (c,d) Temperature dependence of (c) drive and drag currents, $I_{\textrm{drive}}$ and $I_{\textrm{drag}}$, in a drag measurement, and (d) the drag ratio $I_{\textrm{drag}}/I_{\textrm{drive}}$. Measurements are performed at $\nu_{1}=\nu_{2}=1/3$ and $B=28$ T. The gray shaded region indicates the low-temperature regime where both currents vanish. The red shaded region marks the onset of simultaneous $I_{\textrm{drive}}$ and $I_{\textrm{drag}}$, yielding a perfect drag ratio. At higher temperatures, $I_{\textrm{drive}}$ and $I_{\textrm{drag}}$ bifurcate, and the drag ratio deviates from unity. (e) Magnetic field dependence of the counterflow conductance $G_{\text{counter}}$ (log scale) measured at $\nu_{1}=\nu_{2}=1/3$ (black circles), $2/5$ (blue circles), and $3/7$ (red circles).
@@
Quantum statistics form the cornerstone for understanding diverse quantum phases of matter in condensed matter physics. Fermionic statistics fundamentally describe electronic systems at low temperatures, underpinning the physics of Fermi liquids Luttinger1960fermi ; Nozieres2018quantumliquids and the quantum Hall effect Halperin2020fractional . When fermions pair, they form composite particles which are bosons, enabling the emergence of low-temperature ground states characterized by Bose–Einstein condensation. Superconductivity, for instance, results from electrons near Fermi surface undergoing Cooper pairing instabilities Cooper2011BCS . Alternatively, excitonic superfluids represent a distinct class of bosonic condensate, resulting from Coulomb-driven electron-hole pairing, establishing an intriguing phenomenon that is actively pursued in condensed matter research Blatt1962exciton ; Lozovik1975exciton ; Lozovik1976exciton .
@@
The phase space of a quantum Hall bilayer is defined by the Landau level fillings of each graphene layer, $\nu_{1}$ and $\nu_{2}$, which can be independently controlled by adjusting the voltage bias on the top and bottom gate electrodes. Figure 1c presents a schematic phase diagram, where the axes are redefined by the total filling $\nu_{\text{total}}=\nu_{1}+\nu_{2}$ and the layer imbalance $\Delta\nu=\nu_{1}-\nu_{2}$. We focus on the weak-coupling regime dominated by the conventional Jain sequence of FQHE states. Here, red and blue solid lines indicate the expected locations of single-component FQHE states in layers 1 and 2, respectively. Black circles mark a representative set of intersections, where layer 1 is tuned to a Laughlin state at $\nu_{1}=1/3$ and layer 2 is swept across a sequence of Jain states characterized by the effective composite fermion (CF) filling $\nu_{2}^{\ast}=\nu_{2}/(1-2\nu_{2})=N$ with $N\in\mathbb{Z}$ Jain.03 ; Eisenstein1990FQHE . These CFs are constructed using intralayer flux attachment only, distinguishing them from previous reports of two-component FQHE Li2019pairing ; Zhang2025fractionalexciton . In the following, we use $\nu_{i}^{\ast}$ to denote the CF effective filling in layer $i=1,2$.
Across the Jain sequence highlighted in Fig. 1c, the presence of interlayer excitons is revealed through their distinctive transport responses. We employ three distinct electrical transport configurations in bilayer samples patterned into an edgeless Corbino geometry (see Methods for details) Nandi2012exciton ; Eisenstein2014 ; Li2019pairing ; Zeng2023solid ; Zhang2025fractionalexciton . At $B=12$ T, the parallel-flow (PF) geometry exhibits vanishing bulk conductance at $\nu_{2}^{\ast}=N$ (top panel of Fig. 1d), confirming robust charge gaps associated with Jain states. Remarkably, these insulating states in the PF geometry exhibit non-zero conductance in the counterflow configuration (middle panel of Fig. 1d), coinciding with perfect drag responses (bottom panel of Fig. 1d). When a current bias is applied to the drive layer (layer 1), a counter-propagating current of equal magnitude is detected in the drag layer (layer 2), resulting in a unity drag ratio $I_{\text{drag}}/I_{\text{drive}}=1$. This combination of transport signatures is in excellent agreement with prior observations of the (111) state, offering unambiguous evidence for the existence of interlayer excitons across the Jain sequence.
Interestingly, signatures of exciton pairing between a pair of Laughlin state vanishes at $B=28$ T. As shown in Fig. 1e, vanishing conductance is observed in both the PF and counterflow geometries at $\nu_{1}=\nu_{2}=1/3$ (top and middle panels of Fig. 1e). In the drag geometry, both the drive and drag currents approach zero, leading to an ill-defined drag ratio, which is masked with the gray-shaded region in the bottom panel of Fig. 1e.
The Jain sequence of FQHE states, characterized by its hierarchical behavior shown in Fig. 1d-e, is described by well-defined wavefunctions localized within each layer. The absence of interlayer correlation in the ground-state wavefunction raises an intriguing conundrum regarding the origin of interlayer excitons.
The magnetic field dependence of the transport response, as illustrated in Fig. 1d-e, provides an important clue. We propose that exciton pairing occurs between a negatively charged quasiparticle in one layer and a positively charged quasihole in the other. As shown in the inset of Fig. 2a, such quasiparticles and quasiholes are thermally generated across the charge gap. Consequently, their population is governed by the temperature relative to the charge gap. In systems with a large charge gap, such as the situation at $B=28$ T, these excitons are suppressed at low temperatures but are expected to re-emerge as the gap diminishes or as the temperature increases.
This hypothesis is validated by the temperature dependence of the drag response. At $B=28$ T and $\nu_{1}=\nu_{2}=1/3$, drag measurement produces zero $I_{\textrm{drive}}$ and $I_{\textrm{drag}}$ in the low-temperature region, as shown in Fig. 2c. This is consistent with the absence of excitons, and both layers behave like insulators in the bulk. With increasing temperature, quasiparticle excitations are generated across the charge gap, evidenced by non-zero current in both drive and drag layers. Remarkably, the onset of current across two layers recovers a perfect drag ratio, as shown in the red shaded area in Fig. 2d. This observation offers strong support for exciton pairing from thermally excited quasiparticle and quasihole. Upon further increasing $T$, $I_{\textrm{drive}}$ and $I_{\textrm{drag}}$ bifurcate and the drag ratio start to deviate from $1$, indicating that exciton pairing is suppressed at high temperature.
Since the transport response in the parallel flow configuration does not couple to interlayer excitons, the activated behavior in $G_{\text{PF}}$ offers a direct measurement of the underlying charge gap. As illustrated by Fig. S1, the charge gap associated with the Laughlin state at $B=28$T is around $\Delta=40.5$ K, substantially higher than the temperature of current onset in the drag measurement. This discrepancy suggests that exciton pairing lowers the energy cost for creating quasiparticle excitations.
Moreover, the Jain states follow a hallmark hierarchical behavior. As the effective magnetic field decreases at larger effective filling, the charge gap is reduced, making it more preferable to generate quasiparticles and quasiholes. This has a direct impact on the transport response in the counterflow channel. As shown in the bottom panel of Fig. 1e, while the Laughlin state exhibits zero counterflow conductance at $B=28$ T, a perfect drag response is recovered as layer 2 is tuned to Jain states at larger effective fillings.
Along the same vein, Fig. 2e plots counterflow conductance $G_{\text{counter}}$ as a function of $B$ measured at $\nu_{1}^{\ast}=\nu_{2}^{\ast}=1$, $2$, and $3$. At $\nu_{1}^{\ast}=\nu_{2}^{\ast}=1$, exciton population disappears after $B=15$ T, evidenced by vanishing $G_{\text{counter}}$. On the other hand, exciton pairing persists up to $B$ = 30 T at $\nu_{1}^{\ast}=\nu_{2}^{\ast}=2$ and even higher $B$ at $\nu_{1}^{\ast}=\nu_{2}^{\ast}=3$. This comparison establishes a general behavior: at each filling, exciton pairing is suppressed as the charge gap is enhanced with increasing $B$, while at each $B$, excitons become more stable as the charge gap is suppressed at larger effective filling. Combined, our observations strongly support the proposed pairing mechanism between quasiparticles.
Figure 3: Doping neutral anyons into the Laughlin state. (a) Controlling the population of neutral anyons through layer-asymmetric doping, which is achieved by varying Landau level fillings of layer 1 and 2 by ±$\pm\delta$. This generates equal number of quasiparticles and quasiholes across two graphene layers. (b) Parallel flow conductance $G_{\text{PF}}$ (left), drag ratio $I_{\text{drag}}/I_{\text{drive}}$ (middle), and schematic diagram (right) summarizing the transport response across the $\nu_{\text{total}}$-$\Delta\nu$ map around the Laughlin state at $\nu_{1}=\nu_{2}=1/3$ and $B=28$ T. Near $\Delta\nu$=0, $I_{\text{drag}}$ and $I_{\text{drive}}$ both vanish, resulting in ill-defined drag ratio masked by gray-shaded areas. Black arrows in the right panel mark the extend of the robust charge gap along the direction of $\Delta\nu$ and $\nu_{\text{total}}$. (c) $G_{\text{PF}}$, $G_{\text{counter}}$ (top) and $I_{\text{drag}}/I_{\text{drive}}$ (bottom) as a function of $\Delta\nu$ measured at $\nu_{\text{total}}=2/3$. (d) Arrhenius plot of $G_{\text{PF}}$ (black) and $G_{\text{counter}}$ (blue) measured at $\Delta\nu=0.06$. (e) $G_{\text{PF}}$, $G_{\text{counter}}$ as a function of varying $\nu_{1}$ while $\nu_{2}$ is fixed at $1/3$.
Figure 4: Non-Abelian neutral mode at $\nu_{1}=\nu_{2}=3/8$. (a) A detailed view of the $\nu_{\text{total}}$-$\Delta\nu$ phase space, where black and gray lines indicate CF states with two and four flux attachments, respectively. (b) Parallel flow conductance $G_{\text{PF}}$, (c) counterflow conductance $G_{\text{counter}}$, and (d) drag ratio $I_{\text{drag}}/I_{\text{drive}}$ as a function of varying $\nu_{\text{total}}$ measured at $B=31$ T. Vertical arrows highlight different types of charge-neutral anyons emerging at locations marked by red, blue, and green circles in panel (a). (e) $G_{\text{PF}}$ as a function of $\nu_{\text{total}}$ and $B$, revealing a series of $B$-induced transitions in the FQHE states, marked by horizontal black solid lines. (f) $G_{\text{PF}}$ measured across a portion of the phase space encircled by the red dashed line in panel (a). Vanishing $G_{\text{PF}}$ defines incompressible states marked by black and gray circles in panel (a).
In the above mentioned scenario, quasiparticles, quasiholes, and the resulting excitons are all thermal excitations. In the following, we show that such excitons can also be generated as a feature of the low-temperature ground state, which is realized through independently control field-effect doping across two graphene layers.
@@
As illustrated in Fig. 3a, an equal population of quasiparticles and quasiholes can be generated across the two graphene layers by doping the bilayer to $\nu_{1}=1/3+\delta$ and $\nu_{2}=1/3-\delta$. This defines a vertical line in phase space at constant total filling fraction $\nu_{\text{total}}=2/3$, where $\delta$ is directly linked to layer imbalance $\Delta\nu$.
The presence of doping-induced excitons is demonstrated by the transport properties around $\nu_{1}=\nu_{2}=1/3$. Vanishing parallel flow conductance with $G_{\text{PF}}=0$ highlights a region with a fully developed charge gap, represented by dark blue in the left panel of Fig. 3b. While this region spans a range of $0.20$ in the direction of $\Delta\nu$, it covers a much narrower range of only $0.02$ in the direction of $\nu_{\text{total}}$. This directional dependence aligns well with the unique doping requirement for exciton formation. Tuning along other directions, such as varying $\nu_{\text{total}}$ while holding $\Delta\nu$ fixed, leads to mismatched quasiparticle and quasihole populations, resulting in unpaired charges and a lack of drag response.
Within the boundary of vanishing $G_{\text{PF}}$, drag measurements reveal the emergence of exciton pairing. Near $\Delta\nu$ $=0$, the bilayer sample exhibits vanishing conductance in both parallel flow and counterflow measurements, as illustrated in Fig. 3c. As both drive and drag currents are zero, the drag ratio is ill-defined, which is indicated by the gray-shaded area in Fig. 3b and c. As $\Delta\nu$ deviates from zero, the counterflow conductance rises from zero while the parallel flow conductance remains vanishingly small. This leads to a regime characterized by a perfect drag response, represented by dark red in the middle and right panels of panel of Fig. 3b and indicated in the red-shaded area of Fig. 3c. These transport responses point to excitonic pairing, driven by the doping-induced population of quasiparticles and quasiholes.
In the presence of doping-induced excitons, the bulk conductance in the counterflow geometries no longer exhibits activated behavior, as shown in Fig. 3d, while the temperature dependence of $G_{\text{PF}}$ suggests a robust charge gap. This strongly suggests that these doping-induced excitons are intrinsic features of the ground state rather than thermal excitations.
Notably, tuning $\nu_{\text{total}}$ at $\Delta\nu$$=0$ results in a region with a zero drag ratio, outlined by blue dashed lines and highlighted by the blue-shaded area in the middle and right panels of Fig. 3b, respectively. Fig. 3e plots the transport responses as $\nu_{1}$ is detuned from the Laughlin state. The onset of $G_{\text{PF}}$ indicates that the charge gap is suppressed by the presence of quasiparticles in layer 1 (black solid line in Fig. 3e). However, the absence of quasiholes in layer 2 prevents the formation of exciton pairs, thereby suppressing the counterflow conductance, $G_{\text{counter}}$ (blue solid line in Fig. 3e). This results in a non-zero $G_{\text{PF}}$ and a zero $G_{\text{counter}}$, confirming that exciton formation requires doping with varying $\Delta\nu$ while maintaining a constant $\nu_{\text{total}}$.
Having established quasiparticle pairing in the Laughlin state, we now expand our investigation to other types of excitons along the equal density line, defined by $\nu_{1}=\nu_{2}$ and depicted as the horizontal blue line in Fig. 4a. Along this line, transport measurements reveal three distinct types of excitonic states, with their respective positions marked by green, blue, and red circles in Fig. 4a. At $\nu_{1}=\nu_{2}=1/2$ (green circle), the perfect drag response accompanied by vanishing $G_{\text{PF}}$ indicates an excitonic state described by the (111) wavefunction Nandi2012exciton ; Eisenstein2014 ; Zeng2023solid ; Zhang2025fractionalexciton ; Li2017superfluid ; Liu2017superfluid ; Kellogg2003 ; Kellogg2004 ; Tutuc2004counterflow . These excitons are composed of integer-charged constituents that combine to form composite bosons (see inset in Fig. 4b, green frame).
As $\nu_{1}$ and $\nu_{2}$ deviate from half-filling, a series of FQHE states emerge along the Jain sequence, as evidenced by vanishing $G_{\text{PF}}$ at $\nu_{1}^{\ast}=\nu_{2}^{\ast}=2$, 3, 4, and 5 (Fig. 4b). At these fillings, we observe a perfect drag ratio and non-zero counterflow conductance, providing strong support for the presence of interlayer excitons. We argue that the pairing mechanism at these fillings is analogous to that at $\nu_{1}=\nu_{2}=1/3$, where excitons form via interlayer pairing of quasiparticles and quasiholes.

### v3

@@
thanks: These authors contributed equally to this work.thanks: These authors contributed equally to this work.thanks: These authors contributed equally to this work.thanks: Email: jia_\_li@brown.edu
@@
Ron Q. Nguyen1    Naiyuan J. Zhang1    Navketan Khurana-Batra1,2    Sarah Alkidim1    Xiaoxue Liu1    Kenji Watanabe3    Takashi Taniguchi4    D. E. Feldman1,2    J.I.A. Li1,5 1Department of Physics, Brown University, Providence, Rhode Island 02912, USA 2Brown Theoretical Physics Center, Brown University, Providence, Rhode Island 02912, USA 3Research Center for Electronic and Optical Materials, National Institute for Materials Science, 1-1 Namiki, Tsukuba 305-0044, Japan 4Research Center for Materials Nanoarchitectonics, National Institute for Materials Science, 1-1 Namiki, Tsukuba 305-0044, Japan 5Department of Physics, University of Texas at Austin, Austin, TX 78712, USA
(December 13, 2025)
@@
Figure 1: Laughlin states in a quantum Hall graphene bilayer. (a) Schematic of a single-layer Laughlin state stabilized by intralayer Coulomb interactions. At Landau-level filling $\nu=1/3$, its quasiparticle excitations carry a fractional charge of $e/3$. (b) Schematic of two coupled Laughlin states in a quantum Hall bilayer, each stabilized by intralayer interactions. In this configuration, quasiparticles and quasiholes residing in opposite layers interact through weak interlayer Coulomb coupling. (c–d) Counterflow drag measurements as a function of Landau-level filling along the equal-density line $\nu_{1}=\nu_{2}$. Data are acquired from a quantum Hall bilayer device patterned into a Hall-bar geometry. (c) Longitudinal and (d) transverse responses are shown for the drive and drag layers at $B=30$ T and $T=0.3$ K. Insets illustrate the corresponding measurement configurations.
Figure 2: Pairing between quasiparticles and quasiholes of the Laughlin state. (a) Schematic illustrating how the population of bilayer excitons can be tuned by introducing layer-asymmetric doping. This is achieved by varying the Landau-level fillings of layers 1 and 2 by $\pm\delta$, thereby generating equal numbers of quasiparticles and quasiholes in opposite layers. (b) Hall resistance of the drive layer, $R_{xy,\mathrm{drive}}$, (c) longitudinal resistances of the drive and drag layers, $R_{xx,\mathrm{drive}}$ and $R_{xx,\mathrm{drag}}$, and (d) Hall resistance of the drag layer, $R_{xy,\mathrm{drag}}$, as functions of $\Delta\nu$ measured at total filling $\nu_{\mathrm{total}}=2/3$. (e) Parallel-flow conductance $G_{\mathrm{PF}}$, (f) counterflow conductance $G_{\mathrm{counter}}$, and (g) drag current ratio $I_{\mathrm{drag}}/I_{\mathrm{drive}}$ as functions of $\Delta\nu$, measured at the same $\nu_{\mathrm{total}}=2/3$. (h) Map of $G_{\mathrm{PF}}$ and (i) map of $I_{\mathrm{drag}}/I_{\mathrm{drive}}$ across the $\nu_{\mathrm{total}}$$\Delta\nu$ plane near the Laughlin state at $\nu_{1}=\nu_{2}=1/3$ and $B=28$ T. Near $\Delta\nu=0$, both $I_{\mathrm{drag}}$ and $I_{\mathrm{drive}}$ vanish, resulting in an ill-defined drag ratio, indicated by the gray-shaded regions. Panels (b-d) are measured in the Hall-bar-shaped device; panels (e-i) are measured in the Corbino-shaped device.
Figure 3: Exciton binding energy. (a) Arrhenius plot of the parallel-flow current at $\nu_{\mathrm{total}}=2/3$ and $\Delta\nu=0.06$. (b–c) Bulk transport measured at $\nu_{1}=1/3$ as a function of $\nu_{2}$ for (b) $B=28$ T and (c) $B=12$ T. The top, middle, and bottom panels show, respectively, the parallel-flow conductance $G_{\mathrm{PF}}$, counterflow conductance $G_{\mathrm{counter}}$, and drag current ratio $I_{\mathrm{drag}}/I_{\mathrm{drive}}$. (d–e) Temperature dependence of (d) drive and drag currents $I_{\mathrm{drive}}$ and $I_{\mathrm{drag}}$, and (e) the drag ratio $I_{\mathrm{drag}}/I_{\mathrm{drive}}$, measured at $\nu_{1}=\nu_{2}=1/3$ and $B=28$ T. The gray-shaded region marks the low-temperature regime where both currents vanish. The red-shaded region indicates the onset of simultaneous $I_{\mathrm{drive}}$ and $I_{\mathrm{drag}}$, corresponding to a perfect drag ratio. At higher temperatures, $I_{\mathrm{drive}}$ and $I_{\mathrm{drag}}$ bifurcate, and the drag ratio departs from unity. Data in this figure are measured in the Corbino-shaped device.
Figure 4: Exciton pairing at half-filled $\Lambda$-levels. (a) Parallel-flow conductance $G_{\mathrm{PF}}$ as a function of $\nu_{\mathrm{total}}$ and $B$ along the equal-density line, revealing a sequence of magnetic-field–induced transitions in the FQH states, marked by horizontal black lines. In the high-field limit, the FQH states occur at integer and half-integer fillings of the $\Lambda$-levels, indicative of exotic exciton formation. (b) Parallel flow conductance $G_{\mathrm{PF}}$, (c) counterflow conductance $G_{\mathrm{counter}}$, and (d) drag current ratio $I_{\mathrm{drag}}/I_{\mathrm{drive}}$ as functions of $\nu_{\mathrm{total}}$, measured at $B=31$ T. (e) Hall resistance on the drive layer $R_{xy,\mathrm{drive}}$ and longitudinal drag response $R_{xx,\mathrm{drag}}$ as a function of $\nu_{1}=\nu_{2}$ along the equal-density line, measured in a Hall-bar device at $B=25$ T. (f) Longitudinal and Hall drag response, $R_{xx,\mathrm{drag}}$ and $R_{xy,\mathrm{drag}}$, as functions of layer imbalance $\Delta\nu$, measured at $\nu_{1}=\nu_{2}=-3/8$ and $B=25$ T. Panels (a-d) are measured in the Corbino-shaped device; panels (e-f) are measured in the Hall-bar-shaped device.
@@
Quantum statistics form the cornerstone for understanding diverse quantum phases of matter in condensed matter physics. Fermionic statistics fundamentally describe electronic systems at low temperatures, underpinning the physics of Fermi liquids Luttinger1960fermi; Nozieres2018quantumliquids and the quantum Hall effect Halperin2020fractional. When fermions pair, they form composite particles which are bosons, enabling the emergence of low-temperature ground states characterized by Bose–Einstein condensation. Superconductivity, for instance, results from electrons near the Fermi surface undergoing Cooper pairing instabilities Cooper2011BCS. Alternatively, excitonic superfluids represent a distinct class of bosonic condensate, resulting from Coulomb-driven electron-hole pairing, establishing an intriguing phenomenon that is actively pursued in condensed matter research Blatt1962exciton; Lozovik1975exciton; Lozovik1976exciton.
@@
The phase space of a quantum Hall bilayer is defined by the Landau-level fillings of the two graphene layers, $\nu_{1}$ and $\nu_{2}$, which can be independently tuned by adjusting the voltages on the top and bottom gate electrodes. The strength of interlayer coupling is commonly parameterized by the ratio $d/\ell_{B}$, where $d$ is the interlayer separation and $\ell_{B}$ is the magnetic length. In this work, we report measurements from two comparable quantum Hall bilayer samples, each consisting of two monolayer graphene layers separated by an insulating barrier. One device is patterned into a Hall-bar geometry Li2017superfluid; Liu2017superfluid; Liu2022crossover; Eisenstein2014 with $d=5.4$ nm, and the other into a Corbino geometry Zeng2019; Li2019pairing; Zeng2023solid; Zhang2025fractionalexciton; Eisenstein2014; Nandi2012exciton with $d=4.5$ nm. In the lowest Landau level (LL), monolayer graphene is electron–hole symmetric. We therefore focus on the lowest hole-doped LL in each layer, with $\nu_{1}$ and $\nu_{2}$ denoting the corresponding hole filling factors. Further details regarding device fabrication and measurement configurations are provided in the Methods section.
The Laughlin states emerge in the weak-coupling regime around $d/\ell_{B}\sim 1$, where each layer follows the conventional Jain sequence of fractional quantum Hall (FQH) states. Figures 1c–d show counterflow drag measurements from a Hall-bar device as a function of Landau-level filling along the equal-density line, $\nu_{1}=\nu_{2}$. These data reveal a series of Jain states at $\nu_{1}=\nu_{2}=N/(1+2N)$ with $N\in\mathbb{Z}$ Jain.03; Eisenstein1990FQHE. In this regime, the bilayer hosts two decoupled copies of the Laughlin state at $\nu_{1}=\nu_{2}=1/3$. Because each layer possesses a robust charge gap, no free charge carriers are available to support exciton pairing across the layers. Two observations confirm the absence of interlayer correlations: the drive layer exhibits a quantized Hall plateau at $R_{xy,\mathrm{drive}}=h/(\nu e^{2})$, while the drag layer shows negligible signal in both longitudinal and Hall channels.
Exciton pairing between Laughlin states
@@
Starting from the pair of Laughlin states, interlayer excitons can be introduced by asymmetrically doping the graphene bilayer. As illustrated in Fig. 2a, an equal population of quasiparticles and quasiholes is generated across the two layers by tuning the fillings to $\nu_{1}=1/3+\delta$ and $\nu_{2}=1/3-\delta$. This defines a vertical trajectory in the $\nu_{\mathrm{total}}-\Delta\nu$ phase space at constant total filling $\nu_{\mathrm{total}}=2/3$, where the parameter $\delta$ directly quantifies the layer imbalance $\Delta\nu=2\delta$.
The presence of bilayer excitons is revealed by characteristic signatures in the counterflow drag measurements. Figures 2b–d show the drag responses from a Hall-bar device as a function of $\Delta\nu$ at $\nu_{\mathrm{total}}=2/3$. As the bilayer is asymmetrically doped away from $\Delta\nu=0$, the Hall response of the drive layer begins to deviate from the quantized Laughlin plateau around $\Delta\nu\approx-0.08$. Concurrently, pronounced peaks emerge in both the longitudinal and Hall drag resistances, $R_{xx,\mathrm{drag}}$ and $R_{xy,\mathrm{drag}}$, signaling the onset of strong interlayer correlations associated with exciton formation.
Definitive evidence of exciton pairing is obtained from bulk transport measurements using a Corbino-shaped sample. As shown in Figs. 2e–g, the parallel-flow conductance $G_{\mathrm{PF}}$ vanishes within the range $0.1<\Delta\nu<+0.1$, highlighting a regime with a fully developed charge gap. This allows us to identify bilayer excitons through counterflow and drag measurements.
Within this gapped regime, the counterflow conductance also vanishes near $\Delta\nu=0$, as indicated by the gray-shaded region in Figs. 2e–g. Because both the drive and drag currents are zero, the drag ratio is ill-defined in this region. As $\Delta\nu$ is tuned away from zero, the counterflow conductance rises from zero, giving way to a regime where finite counterflow conductance coexists with vanishing parallel-flow conductance, denoted by the red-shaded area in Figs. 2e–g. Most remarkably, drag measurements reveal a perfect drag response in this regime, with a drag ratio of unity ($I_{\mathrm{drag}}/I_{\mathrm{drive}}=1$). Together, these transport responses provide compelling evidence for the formation of interlayer excitons. Remarkably, in this regime the only charge-carrying entities are excitons, which sustain counterflowing currents while no free charge transport occurs in the parallel-flow channel.
Figures 2h–i present bulk transport data around $\nu_{1}=\nu_{2}=1/3$. The regime with a robust charge gap is identified by vanishing parallel-flow conductance, represented by dark blue in Fig. 2h. While this region extends over a range of approximately $0.20$ along the $\Delta\nu$ axis, it spans only about $0.02$ along the $\nu_{\mathrm{total}}$ axis. This pronounced anisotropy suggests that the charge gap is stabilized by exciton formation along the direction of varying $\Delta\nu$ at constant $\nu_{\mathrm{total}}$. Indeed, a perfect drag response—represented by red in Fig. 2i—is observed near the tips of the elongated charge-gap region, where finite $\Delta\nu$ is expected to induce interlayer excitons. In contrast, varying $\nu_{\mathrm{total}}$ at $\Delta\nu=0$ closes the charge gap and restores finite parallel-flow conductance.
Thermal excitation and exciton binding energy
The formation of excitons is associated with a characteristic binding energy, which is characterized in the following. Figure 3a shows the Arrhenius plot of the parallel-flow conductance measured at $\nu_{\mathrm{total}}=2/3$ and $\Delta\nu=0.06$, where a perfect drag response is observed at low temperature. The activated behavior reflects the energy required to generate free charge carriers from a ground state containing excitons. Within this interpretation, the energy gap extracted from the Arrhenius plot provides a lower bound for the exciton binding energy.
Figures 3b–c present transport measurements obtained by sweeping $\nu_{2}$ while holding $\nu_{1}$ fixed at $1/3$. At $B=28$ T, the parallel-flow conductance vanishes at $\nu_{2}=N/(1+2N)$ (top panel of Fig. 3b), consistent with the robust charge gaps of the Jain sequence. Owing to the large charge gap at $\nu_{1}=\nu_{2}=1/3$, excitons are absent, as indicated by the vanishing counterflow conductance and the ill-defined drag ratio.
Remarkably, exciton pairing at $\nu_{1}=\nu_{2}=1/3$ re-emerges at $B=12$ T. As shown in Fig. 3c, all Jain states—including the Laughlin state at $\nu_{1}=\nu_{2}=1/3$—display finite counterflow conductance (middle panel), accompanied by perfect drag responses (bottom panel). Compared to the high-field measurement at $B=28$ T, the reduced magnetic field weakens the charge gap at $\nu_{1}=\nu_{2}=1/3$, thereby allowing quasiparticle excitations at low temperature and facilitating exciton formation. A corresponding enhancement of the drag response at $\nu_{1}=\nu_{2}=1/3$ is also observed in the Hall-bar device at $B=12$ T, as illustrated in Fig. M2d.
Interestingly, the Laughlin state at $\nu_{1}=\nu_{2}=1/3$ exhibits a charge gap of $\Delta=12$ K at $B=12$ T, which is substantially higher than the temperature of $\sim 20$ mK where the perfect drag response is observed. This observation strongly suggests that, while the energy cost of generating individual quasiparticle excitations from the Laughlin state is $\Delta=12$ K, the pairing between positively and negatively charged constituents dramatically reduces the effective energy required to form excitons. Hence, $\Delta=12$ K provides an additional estimate for the lower bound of the exciton binding energy at $B=12$ T. Moreover, our findings suggest that exciton pairing between Laughlin states occurs by binding quasiparticle and quasihole excitations across graphene layers.
This hypothesis is validated by the temperature dependence of the drag response at $B=28$ T and $\nu_{1}=\nu_{2}=1/3$, where thermal activation induces a perfect drag response emerging from an insulating ground state. As shown in Fig. 3d, the absence of excitons in the low-temperature regime results in vanishing $I_{\mathrm{drive}}$ and $I_{\mathrm{drag}}$, indicating that both graphene layers behave as insulators. With increasing temperature, both the drive and drag currents begin to rise around $T\approx 2$ K, which is far below the charge gap extracted from PF conductance (see Fig. 3a). Remarkably, this reveals a narrow temperature window of perfect drag response, highlighted by the red-shaded region in Fig. 3e. This observation provides compelling evidence for exciton pairing between thermally excited quasiparticles and quasiholes. Upon further increasing temperature, $I_{\mathrm{drive}}$ and $I_{\mathrm{drag}}$ bifurcate, and the drag ratio deviates from unity, indicating that exciton pairing is suppressed at elevated temperatures.